OpenStudy (anonymous):

A red die and a green die have been rolled. What is the probability of the sum is 9

OpenStudy (nikvist):

sum 9: (3,6) , (4,5) , (5,4) , (6,3) -> total_9=4 any sum: 6*6=36 probability=4/36=1/9

OpenStudy (anonymous):

thanks